# STaRK: Benchmarking LLM Retrieval on Textual and Relational Knowledge Bases

### v1

@@ Abstract
Answering real-world user queries, such as product search, often requires accurate retrieval of information from semi-structured knowledge bases or databases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information. However, previous works have mostly studied textual and relational retrieval tasks as separate topics. To address the gap, we develop STaRK, a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases. We design a novel pipeline to synthesize natural and realistic user queries that integrate diverse relational information and complex textual properties, as well as their ground-truth answers. Moreover, we rigorously conduct human evaluation to validate the quality of our benchmark, which covers a variety of practical applications, including product recommendations, academic paper searches, and precision medicine inquiries. Our benchmark serves as a comprehensive testbed for evaluating the performance of retrieval systems, with an emphasis on retrieval approaches driven by large language models (LLMs). Our experiments suggest that the STaRK datasets present significant challenges to the current retrieval and LLM systems, indicating the demand for building more capable retrieval systems that can handle both textual and relational aspects.
Figure 1: STaRK features queries on semi-structured knowledge bases (SKBs) with textual and relational knowledge, with node entities as ground truth answers. It simulates user interactions with a SKB to evaluate LLM retrieval systems’ performance in providing accurate responses.
Example query Title of ground truth node(s)
@@
resist easy tearing. Any recommendations? <Dart World Broken Glass Flight> (12 more)
@@ 1 Introduction
Natural-language queries are the primary forms in human society for acquiring information, involving diverse knowledge in the real world [15, 23, 19]. For example, users on e-commerce web search can express complex information needs by combining free-form elements or constraints, such as “Can you help me find a push-along tricycle from Radio Flyer that’s both fun and safe for my kid?”. Here, the underlying knowledge can be represented in semi-structured formats [30, 41, 33], referred to as semi-structured knowledge bases (SKBs), which integrate unstructured data, such as textual descriptions and natural language expressions (e.g., descriptions about a tricycle), with structured data, like entity interactions on knowledge graph or database (e.g., a tricycle with “belongs to” relation with brand Radio Flyer). This allows semi-structured knowledge bases to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [13], social media [27], and biomedicine [6, 16]. In this work, we focus on semi-structured knowledge bases (SKBs) with texts and relation information, which are two prevalent modalities on the existing knowledge bases.
In this context, as shown in Figure 1, information retrieval serves as a predominant step by identifying relevant data from the semi-structured knowledge bases, such as specific tricycles that match the request, which are then processed to generate accurate answers to user queries [35]. The retrieval accuracy on semi-structured knowledge bases is crucial for enhancing user experience, supporting informed decision-making, and preventing hallucination.
@@
Limitations of existing works. However, previous works and benchmarks have primarily focused on either purely textual queries on unstructured knowledge [18, 12, 22, 25, 44] or structured SQL or knowledge graph queries [2, 48, 46, 37], which are inadequate to study the complexities of retrieval tasks involving semi-structured knowledge bases. Another line of works [10, 21, 46, 42, 43, 38] have explored the middle ground via multi-hop reasoning within single or across multiple documents. However, they are either limited in the span of knowledge or lack of diverse textual properties and explicit relational information. Recently, large language models (LLMs) have demonstrated significant potential on information retrieval tasks [52, 12, 26, 34]. Nevertheless, it remains an open question how effectively LLMs can be applied to the specific challenge of retrieval from SKBs. It is also important to note that existing works focus mainly on general knowledge, e.g., from wikipedia. In fact, the knowledge may also come from private sources, necessitating information retrieval systems adaptable to private semi-structured knowledge bases. Therefore, there is a gap in our understanding of how current LLM retrieval systems handle the complex interplay between textual and relational requirements in queries that optionally involve private knowledge.
@@
Challenges. To address this gap, our goal is to develop a benchmark for LLM retrieval on SKBs. However, the lack of real queries on SKBs has become an obstacle. Although retrieval datasets with synthetic queries is a valuable supplement, accurately simulating user queries on SKBs is particularly difficult. This difficulty arises from the interdependence of textual and relational information, which can lead to challenges in precisely filtering through millions of candidates to construct the ground truth answers. Additionally, ensuring that queries are relevant to real-world applications and resembles real-world scenarios adds further complexity to the benchmarking process.
@@
Our contribution: STaRK. As shown in Figure 1, we present a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases (STaRK). We propose a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources. We validate the quality of queries in our benchmark through detailed analysis and human evaluation, focusing on their naturalness, diversity, and practicality. With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and providing insights for future advancement. We point out the key features of STaRK as follows:
@@
  • Natural-sounding queries on semi-structured knowledge: As illustrated in Table 1, the queries in our benchmark are crafted to incorporate rich relational information and complex textual properties. Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with extra contexts.
@@
  • Diverse domains: Our benchmark spans a wide range of content and domains, reflecting the diverse nature of real-world information needs. It covers areas on product recommendation, academic paper search, and precision medicine inquiry. Therefore, STaRK provides a comprehensive evaluation of retrieval systems across various contexts and knowledge domains.
Moreover, we perform extensive experiments on each benchmark dataset for the prevailing retrieval systems. We highlight research challenges especially on the capability to handle textual and relational requirements and the latency of retrieval systems on large-scale SKBs that involve million-scale entities or relations. Finally, we provide insights into future directions to build more capable retrieval systems. The benchmark datsets and code are available on https://github.com/snap-stanford/stark.
2 Related Work
Unstructured QA Datasets. The line of work explores retrieving answers from a single document [32] or multiple document sources [40, 45, 21, 42, 10]. For example, SQuAD [32] focuses on extracting answers from a single document, testing model’s ability to understand and retrieve information within a specific context. For multi-document retrieval, HotpotQA [45] demands reasoning over multiple documents and includes sentence-level supporting facts, enhancing the complexity of the retrieval task. TriviaQA [21] combines questions from trivia games with evidence documents, evaluating both retrieval and reasoning over multiple sources. Moreover, some works [29, 24] leverage generated results by search engines as ground truth answers or information sources for retrieval. However, unstructured QA datasets or knowledge bases often lack the depth of relational reasoning commonly required in answering complex user queries.
Structured QA Datasets. Within this category, the datasets challenge models to retrieve answer from structured knowledge bases, either knowledge graphs [3, 49, 39, 4, 5, 11, 14] or tabular data [51, 50]. For example, ComplexWebQuestions [39] challenge model’s abilities to decompose complex queries and interpret queries using entities and relationships defined in knowledge graphs. GraphQA [14] textualizes graphs by describing node attributes and edges with natural language, offering language model a way to handle KBQA. For datasets with tabular data, WikiSQL [51] and Spider [50] both address the task of converting natural language questions to SQL queries, with WikiSQL concentrating on single-table databases and Spider extending to complex multi-tables. While these datasets primarily focus on relational information, the lack of textual information limits questions within predefined relationships and entities, which constrains the breadth of available information, resulting in less comprehensive and nuanced responses compared to the extensive insights drawn from abundant textual data.
Semi-Structured QA Datasets. Another line of work focuses on semi-structured QA with tabular and textual data. WikiTableQuestions [31] emphasizes comprehension of table structure and textual content, reflecting the semi-structured nature of many real-world datasets. TabFact [7] and HybridQA [8] combine textual and tabular data, requiring models to validate statements or answer questions using both data types. TabMCQ [20] introduces multiple-choice questions based on tables, adding complexity in understanding and reasoning over semi-structured data. However, different from this work, these existing datasets mainly focus on tables as structured sources, which may not incorporate rich relational information naturally exists between entities. Additionally, previous attempts to combine text and tabular data through external web links or additional text columns often result in data structures that are hard to navigate. Therefore, there is an emerging need to benchmark retrieval tasks on textual and relational knowledge base, offering more insights beyond the scope of structured tabular sources.
@@ 3 Benchmarking Retrieval Tasks over Textual and Relational Knowledge
STaRK-Amazon 2 3 3.0 1,032,407 3,886,603 592,067,882
STaRK-MAG 4 4 10.6 1,872,968 19,919,698 212,602,571
STaRK-Prime 10 18 62.6 129,375 8,100,498 31,844,769
We develop three large-scale retrieval datasets on three semi-structure knowledge bases (SKBs). In Section 3.1, we introduce the schema and scale of the SKBs, which integrate entities with rich relational and textual information. In Section 3.2, we introduce the retrieval datasets and the key characteristics of the queries. Then, we highlight our novel dataset construction pipeline in Section 3.3, which simulates user queries involving textual and relational knowledge and automatically generates the ground truth answers in the retrieval datasets. Finally, we conduct data distribution analysis and human evaluation on the queries in Section 3.4.
3.1 Semi-structured Knowledge Bases
@@
Amazon Semi-structured Knowledge Base. This knowledge base is derived from the Sports and Outdoors category from Amazon Product Reviews [13] and Amazon Question and Answer Data [28]. The knowledge base features two entity types: product and brand, and three types of relational information: also_bought, also_viewed between product entities, and has_brand relation between product and brand entities. In total, it comprises around 1.03M entities (product entities: 0.96M, brand entities: 0.07M) and 3.9M relations (also_bought: 1.7M, also_viewed: 1.3M, has_a_brand: 0.9M). We obtain the textual information by combining the meta data from Amazon Product Reviews with the customer Q&A records from Amazon Question and Answer Data. This provides a rich amount of textual data, including product titles, descriptions, prices, customer reviews, and related customer Q&A for each product. For brand entities, we extract brand titles as the textual attribute. Especially, Amazon SKB features an extensive textual data, which is largely contributed from customer reviews and Q&A.
@@
MAG Semi-structured Knowledge Base. This knowledge base is constructed from obgn-MAG [17], obgn-papers100M [17], and Microsoft Academic Graph (version 2019-03-22) [41, 36]. The knowledge base contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 9K, field_of_study: 0.06M) and 20M relations under four relation types (author_writes_paper: 7.9M, paper_has_field_of_study: 7.2M, paper_cites_paper: 4.9M, author_affiliated_with_institution: 1.0M). We construct the knowledge base by selecting papers1 shared between obgn-MAG and obgn-papers100M. The relational structure comes from the extracted subgraph in obgn-mag based on the selected papers. The textual information such as titles and abstracts is sourced from obgn-papers100M. Additionally, we augment the textual data by integrating details from the Microsoft Academic Graph database, providing extra information like paper venue, author and institution names. This SKB demonstrates a large number of entities and relations associate with paper nodes, especially on citation and authorship relation types.
@@
Prime Semi-structured Knowledge Base. We leverage the existing knowledge graph PrimeKG [6] which contains ten entity types including disease, drug, gene/protein, and eighteen relation types, such as associated_with, synergistic_interaction, indication, expression_present. The entity count in our knowledge base is approximately 129K, with around 8M relations. The details on entity numbers and relation tuples are available in Appendix A.2. Compared to the Amazon and MAG SKBs, Prime SKB is denser and features a greater variety of relation types. While PrimeKG already provides text information on disease and drug entities, including drug mechanisms and disease descriptions, we additionally integrate the textual details from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event.
We include the comprehensive public sources we used to construct the SKBs in Table 8 of Appendix A.
@@ 3.2 Retrieval Tasks on Semi-structured Knowledge Bases
Table 3: Dataset statistics on STaRK.
#queries #queries w/ avg. #answers train / val / test
multiple answers
STaRK-Amazon 9,100 7,082 17.99 0.65 / 0.17 / 0.18
@@
We develop three novel retrieval datasets. Specifically, the task is to retrieve node entities over the SKBs given any input queries. To highlight, the queries feature the interplay and fusion between relational and textual knowledge. Moreover, to elevate their applicability in practical scenarios, these queries are designed to mimic real-life query patterns in terms of the natural-sounding property and flexible formats. Each query may have single answer or multiple answers, formulated as a subset of entities from the knowledge base, which requires the model to accurately identify the ground truth entities out of potentially millions in response to each query. In Table 1, we demonstrate example queries from these benchmarks. We detail our retrieval benchmarks’ scale in Table 3 and introduce the specifics of each retrieval dataset as follows:
@@
STaRK-Amazon. This dataset comprises 9,100 customer-focused queries aimed at product searches, with a notable 68% of these queries yielding multiple answers. The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style. Additionally, it accentuates relational aspects, including brand and product connections (e.g., complementary or substitute items). The queries are framed in detailed, conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
@@
STaRK-MAG. This dataset comprises 13,323 paper queries, with roughly half yielding multiple answers. Beyond the single-hop relational requirements, STaRK-MAG also emphasizes the fusion between the textual requirements with multi-hop queries. For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure. We designed three single-hop and four multi-hop relational query templates, cf. Appendix A.1. The textual aspects of the queries diversifies, focusing on different elements such as the paper’s topic, methodology, or the advancements it introduces, sourced mainly from the abstracts.
@@
STaRK-Prime. This dataset, comprising 11,204 queries across all ten entity types, incorporates single-hop and multi-hop relational information similar to STaRK-MAG. We developed 28 multi-hop query templates, detailed in Appendix A.2, to cover various relation types and ensure their practical relevance. For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments. For entities like drug, disease, gene/protein, and pathway, the queries are a hybrid of relational requirements and textual requirements drawn from documents. A notable feature is the deliberate simulation of three distinct roles – medical scientist, doctor, and patient – for certain queries related to drugs and diseases. This is intended to diversify the language used across various user types and to evaluate the robustness of the retrieval system under varied linguistic scenarios. For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information in the knowledge base.
@@
We divide each dataset into training, validation, and testing subsets, with the ratios detailed in Table 3. For the STaRK-Amazon dataset, we randomly allocate a subset of queries with 20 or fewer answers to the validation and testing sets, while the remaining queries are assigned to the training set. For the STaRK-MAG and STaRK-Prime, we use random splits to divide the data.
3.3 Benchmark Construction

### v2

@@ Abstract
Answering real-world complex queries, such as complex product search, often requires accurate retrieval from semi-structured knowledge bases that involve blend of unstructured (e.g., textual descriptions of products) and structured (e.g., entity relations of products) information. However, previous works have mostly studied textual and relational retrieval tasks as separate topics. To address the gap, we develop STaRK, a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases. Our benchmark covers three domains/datasets: product search, academic paper search, and queries in precision medicine. We design a novel pipeline to synthesize realistic user queries that integrate diverse relational information and complex textual properties, together with their ground-truth answers (items). We conduct rigorous human evaluation to validate the quality of our synthesized queries. We further enhance the benchmark with high-quality human-generated queries to provide an authentic reference. STaRK serves as a comprehensive testbed for evaluating the performance of retrieval systems driven by large language models (LLMs). Our experiments suggest that STaRK presents significant challenges to the current retrieval and LLM systems, indicating the demand for building more capable retrieval systems. The benchmark data and code are available on https://github.com/snap-stanford/stark.
Figure 1: STaRK features queries on Semi-structured Knowledge Base (SKB) with textual and relational knowledge, with node entities as ground-truth answers. STaRK consists of synthesized queries simulating user interactions with a SKB and human-generated queries which provide an authentic reference. It evaluates LLM retrieval systems’ performance in providing accurate responses.
Example query Title of ground-truth items(s)
@@
resist easy tearing. Any recommendations? <Dart World Broken Glass Flight> (+12 more)
@@ 1 Introduction
Natural-language queries are the primary form of how humans acquire information [15, 24, 19]. For example, users on e-commerce sites wish to express complex information needs by combining free-form elements and constraints, such as “Can you help me find a push-along tricycle from Radio Flyer that’s both fun and safe for my kid? in product search. Medical scientists may ask questions like “What disease is associated with the PNPLA8 gene and presents with hypotonia as a symptom?. Answering such queries is crucial for enhancing user experience, supporting informed decision-making, and preventing hallucination.
@@
To answer such queries the underlying knowledge can be represented in semi-structured knowledge bases (SKBs) [31, 43, 34], which integrate unstructured data, such as textual descriptions and natural language expressions (e.g., description of the tricycle), with structured data, like entity interactions on knowledge graph (e.g., a tricycle “brand” is Radio Flyer). This allows the SKBs to represent comprehensive knowledge in specific applications, making them indispensable in domains such as e-commerce [13], social media [28], and precision medicine [6, 16, 21].
@@
Limitations of prior works. Prior works and benchmarks have primarily focused on either purely textual queries on unstructured knowledge [18, 12, 23, 26, 46] or structured SQL or knowledge graph queries [2, 50, 48, 38], which are inadequate to study the complexities of retrieval tasks involving SKBs. Another line of works [10, 22, 48, 44, 45, 39] have explored the middle ground via multi-hop reasoning within single or across multiple documents. However, they are either limited in the span of knowledge or lack of diverse textual properties and explicit relational information. Recently, large language models (LLMs) have demonstrated significant potential on information retrieval tasks [54, 12, 27, 36]. Nevertheless, it remains an open question how effectively LLMs can be applied to the specific challenge of retrieval from SKBs. It is also important to note that existing works focus mainly on general knowledge, e.g., from Wikipedia. In fact, the knowledge may commonly come from private sources, requiring information retrieval systems to operate on private SKBs. Therefore, there is a gap in our understanding of how current LLM retrieval systems handle the complex interplay between textual and relational requirements in queries that optionally involve private knowledge.
@@
Present work. To address this gap, we present a large-scale Semi-structure retrieval benchmark on Textual and Relational Knowledge Bases (STaRK) (Figure 1). The key technical challenge that we solve is how to accurately simulate user queries on SKBs. This difficulty arises from the interdependence of textual and relational information, which can lead to challenges in precisely filtering through millions of candidates to construct the ground-truth answers. Additionally, ensuring that queries are relevant to real-world applications and resembles real-world scenarios adds further complexity to the benchmarking process.
@@
We develop a novel pipeline that simulates user queries and constructs precise ground truth answers using three SKBs built from extensive texts and millions of entity relations from public sources. We validate the quality of queries in our benchmark through detailed analysis and human evaluation, focusing on their naturalness, diversity, and practicality. Furthermore, we incorporate 274 human-generated queries into our benchmark to compare with synthesized queries and enrich the testing scenarios. With STaRK, we delve deeper into retrieval tasks on SKBs, evaluate the capability of current retrieval systems, and provide insights for future advancement. Key features of STaRK are:
@@
  • Natural-sounding queries on SKBs (Table 1): Queries in our benchmark are crafted to incorporate rich relational information and complex textual properties. Additionally, these queries closely mirror the types of questions users would naturally ask in real-life scenarios, e.g., with flexible query formats and possibly with extra contexts.
@@
  • Diverse domains: Our benchmark spans a wide range of domains, reflecting the diverse nature of real-world information needs. It covers three different knowledge bases1footnotetext: 1 We provide an interactive interface for exploring the SKBs at https://huggingface.co/spaces/snap-stanford/SKB-Explorer with applications on product recommendation, academic paper search, and precision medicine inquiry. Therefore, STaRK provides a comprehensive evaluation of retrieval systems across various contexts and knowledge domains.
We also perform extensive experiments on each benchmark dataset for the prevailing retrieval systems. We highlight research challenges especially on the capability to handle textual and relational requirements and the latency of retrieval systems on large-scale SKBs that involve million-scale entities or relations. Finally, we provide insights into future directions to build more capable retrieval systems over private knowledge bases.
@@ 2 Benchmarking Retrieval Tasks over Textual and Relational Knowledge
STaRK-Amazon 4 4 18.2 1,035,542 9,443,802 592,067,882
STaRK-MAG 4 4 43.5 1,872,968 39,802,116 212,602,571
STaRK-Prime 10 18 125.2 129,375 8,100,498 31,844,769
2.1 Problem Definition
We are given a Semi-Structured Knowledge Base (SKB), which consists of a knowledge graph $G$ and a collection of free text documents $D$. Formally, let $G=(V,E)$ be the knowledge graph, where $V$ is the set of nodes and $E\subseteq V\times V$ is the set of edges representing relationships between nodes. $D=\bigcup_{i\in V}D_{i}$ be the collection of free-form text documents associated with the nodes, where $D_{i}$ is the set of documents associated with node $i$. For example, the product knowledge graph in e-commerce can capture relationships between products and brands/colors/categories, and the corresponding text documents include product descriptions, reviews, etc..
We define the tasks on our benchmark datasets as follows: Given the knowledge graph $G=(V,E)$, a collection of free text documents $D$, and a query $Q$, the output is a set of nodes $A\subseteq V$ such that for each node $i\in A$, it satisfies the relational requirements imposed by the structure of $G$ as specified in $Q$, and the associated documents $D_{i}$ satisfy the textual requirements specified in $Q$.
2.2 Semi-structured Knowledge Bases (SKBs)
@@
Amazon Semi-structured Knowledge Base. This knowledge base is derived from the Sports and Outdoors category from Amazon Product Reviews [13] and Amazon Question and Answer Data [29]. The knowledge base features four entity types: product, brand, color, and category, and four types of relational information: also_bought, also_viewed between product entities, has_brand, has_color, has_category relations between product and brand, color, and category entities, respectively. In total, it comprises around 1.0M entities (product entities: 0.9M, brand entities: 0.1M, category entities: 1.4k, color entities: 1.7k) and 9.4M relations (also_bought: 2.8M, also_viewed: 1.9M, has_brand: 1.7M, has_category: 2.3M, has_color: 0.6M). We obtain the textual information by combining the meta data from Amazon Product Reviews with the customer Q&A records from Amazon Question and Answer Data. This provides a rich amount of textual data, including product titles, descriptions, prices, customer reviews, and related customer Q&A for each product. For other entities, we extract their names or titles as the textual attribute. Amazon SKB features an extensive textual data, which is largely contributed from customer reviews and Q&A.
@@
MAG Semi-structured Knowledge Base. This knowledge base is constructed from obgn-MAG [17], obgn-papers100M [17], and Microsoft Academic Graph (version 2019-03-22) [43, 37]. The knowledge base contains around 1.9M entities under four entity types (author: 1.1M, paper: 0.7M, institution: 8.7K, field_of_study: 59.5k) and 39.8M relations under four relation types (author_writes_paper: 13.5M, paper_has_field_of_study: 14.5M, paper_cites_paper: 9.7M, author_affiliated_with_institution: 2.0M). We construct the knowledge base by selecting papers1 shared between obgn-MAG and obgn-papers100M. The relational structure comes from the extracted subgraph in obgn-mag based on the selected papers. The textual information such as titles and abstracts is sourced from obgn-papers100M. Additionally, we augment the textual data by integrating details from the Microsoft Academic Graph database, providing extra information like paper venue, author and institution names. This SKB demonstrates a large number of entities and relations associate with paper nodes, especially on citation and authorship relation types.
@@
Prime Semi-structured Knowledge Base. We leverage the existing knowledge graph PrimeKG [6] which contains ten entity types including disease, drug, gene/protein, and eighteen relation types, such as associated_with, synergistic_interaction, indication, expression_present. The entity count in our knowledge base is approximately 129.3K, with around 8.1M relations. The details on entity numbers and relation tuples are available in Appendix B.3. Compared to the Amazon and MAG SKBs, Prime SKB is denser and features a greater variety of relation types. While PrimeKG already provides text information on disease and drug entities, including drug mechanisms and disease descriptions, we additionally integrate the textual details from multiple databases for gene/protein and pathway entities such as genomic position, gene activity summary and pathway orthologous event. We include the public sources for SKBs construction in Table 10 (Appendix B).
@@ 2.3 Retrieval Tasks on Semi-structured Knowledge Bases
Table 3: Statistics on the STaRK benchmark datasets.
#queries #queries w/ average train / val / test
multiple answers #answers
Synthesized (Sec 2.4, 2.5) STaRK-Amazon 9,100 7,082 17.99 0.65 / 0.17 / 0.18
@@
Human-generated (Sec 2.6) STaRK-Amazon 81 64 19.50 For testing only
STaRK-MAG 84 34 3.26
STaRK-Prime 109 43 2.61
@@
We develop three novel retrieval-based question-answering datasets. To highlight, the queries feature the synthesization between relational and textual knowledge. Moreover, to elevate their applicability in practical scenarios, these queries are designed to mimic real-life query patterns in terms of the natural-sounding property and flexible formats. We detail our retrieval benchmarks’ scale in Table 3 and introduce the specifics of each retrieval dataset as follows:
@@
STaRK-Amazon. This dataset comprises synthesized 9,100 customer-focused queries along with 81 human-generated queries for additional evaluation. The task aims at product searches, with a notable 68% of the synthesized queries yielding multiple answers. The dataset prioritizes customer-oriented criteria, highlighting textual elements such as product quality, functionality, and style. Additionally, it accentuates relational aspects, including brand, category, color and product connections (e.g., complementary or substitute items). The queries are framed in detailed, conversation-like formats, enriching the context and enhancing the dataset’s relevance to real-world scenarios.
@@
STaRK-MAG. This dataset comprises 13,323 synthesized and 84 human-generated queries targeting at academic paper search. Beyond the single-hop relational requirements on STaRK-Amazon, STaRK-MAG also emphasizes the fusion between the textual requirements with multi-hop queries. For example, “Are there any papers from King’s College London” highlights the metapath (institution $\rightarrow$ author $\rightarrow$ paper) on the relational structure. We designed three single-hop and four multi-hop relational query templates, cf. Appendix B.2. The textual aspects of the queries diversifies, focusing on different elements such as the paper’s topic, methodology, or the advancements it introduces, sourced mainly from the abstracts.
@@
STaRK-Prime. This dataset, comprising 11,204 synthesized queries and 109 human-generated queries across all ten entity types, incorporates single-hop and multi-hop relational information similar to STaRK-MAG. For synthesized queries, we developed 28 multi-hop query templates, detailed in Appendix B.3, to cover various relation types and ensure their practical relevance. For instance, the query “What is the drug that targets the genes or proteins expressed in <anatomy>?” serves applications in precision medicine and pharmacogenomics, aiding researchers and healthcare professionals in identifying drugs that act on genes or proteins associated with specific anatomical areas and enabling more targeted treatments. For entities like drug, disease, gene/protein, and pathway, the queries are a hybrid of relational requirements and textual requirements drawn from documents. A notable feature is the deliberate simulation of three distinct roles – medical scientist, doctor, and patient – for certain queries related to drugs and diseases. This is intended to diversify the language used across various user types and to evaluate the robustness of the retrieval system under varied linguistic scenarios. For entities such as effect/phenotype, the queries rely solely on relational data due to limited textual information in the knowledge base.
@@
We divide the synthesized queries on each dataset into training, validation, and testing subsets, with the ratios detailed in Table 3, while using the human-generated queries as additional sets for evaluation. For the STaRK-Amazon dataset, we randomly allocate a subset of queries with 20 or fewer answers to the validation and testing sets, while the remaining queries are assigned to the training set. For the STaRK-MAG and STaRK-Prime, we use random splits to divide the data.
2.4 Benchmark Construction: Synthesized Queries